Laika is pushing a box up an incline of 30?. Laika

Laika is pushing a box up an incline of 30?. Laika exerts a force of 20N in the positive x direction (not striaght up the incline, but into the incline, as shown). Given that the box started at rest (initial velocity = 0m/s), has a mass of 2kg, and is accelerating up the incline at a constant rate of 0.5m/s2 for 4 seconds, what is the net Work done on the box? what is the work done by gravity?
Answer #1

speed of the block after 4 seconds, v = u + a*t

= 0 + 0.5*4

= 2 m/s

a) net workdone on the block = change in kinetic energy of the block

= 0.5*m*(v^2 - u^2)

= 0.5*2*(2^2 - 0^2)

= 4 J

b) distance travelled along the incline in 4 seonds,
d = u*t + 0.5*a*t^2

= 0 + 0.5*0.5*4^2

= 4 m

workdone by gravity = -m*g*h

= -m*g*d*sin(30)

= -2*9.8*4*sin(30)

= -39.2 J